# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ 13. Open Research Challenges
Pre-processing legal texts is challenging due to the distinct nature of legal documents. Existing legal corpora are often contained of raw texts that require extensive cleaning and transformation to become suitable for ML models. Additionally, legal documents can include complex nested structures, such as clauses within clauses, and cross-references to other legal cases, statutes, or provisions, making it difficult to break them into coherent units for analysis. These challenges of legal documents make it impractical to directly fine-tune LMs on these raw datasets without substantial pre-processing. This requirement complicates the use of large legal datasets, making them hard to convert into formats that NLP models can readily process and learn from. Without addressing these specific complexities, fine-tuning LMs on raw legal data becomes impractical, limiting the performance of legal NLP applications.
@@
The use of RLHF within the legal domain is notably scarce. Currently, there is only one peer-reviewed work (Nguyen et al., 2021) available that explores this approach. This indicates an opportunity for research and development in this area, as RLHF could potentially enhance NLP’s capability to learn and make decisions based on complex legal data under human guidance. Further exploration into this method could lead to more responsive and adaptable legal NLP systems. However, due to the complex nature of legal reasoning and the need for accurate legal knowledge in human feedback phase, RLHF’s integration into legal NLP poses some challenges. Therefore, on the human feedback side, legal experts such as lawyers and judges must provide guidance to ensure the AI accurately interprets and applies complex legal concepts.
@@
There is a noticeable gap in the research across various areas of the legal domain, including Intellectual Property, Criminal Law, Banking Law, Family Law, and Human Rights Law. These fields have seen limited exploration across all legal NLP tasks, such as LQA and other applications. Expanding research into these areas is essential for developing comprehensive legal automated systems that can provide tailored solutions and insights highly relevant to these sectors of law.
@@
Research into SLMs specific to the legal domain is notably absent. Addressing this gap could lead to more efficient, resource-conscious solutions that still maintain high performance in legal text processing and analysis. The development of SLMs tailored for legal applications could revolutionise the accessibility and scalability of legal NLP tools.
@@
Domain-specific efficient fine-tuning within the legal field is an underexplored area, with only two known studies (Louis et al., 2024; Li et al., 2022) addressing it. Legal texts consist of complex structures and specialised words that standard LMs may not capture without significant adaptation. Additionally, the legal domain cover a vast array of document types, such as case law, statutes, and contracts, each requiring tailored approaches for effective model application. This diversity makes it imperative to develop fine-tuning strategies that do not only adapt a model generally but rather tailor it to understand the differences between these document types. The majority of existing approaches involve fine-tuning the entire model, which can be resource-intensive. More focused research could enable fine-tuning of legal LMs using fewer resources, enhancing the efficiency of deploying these models in practice.
@@
Complex legal logical reasoning remains a challenge in LJP, particularly in predicting prison terms. Current state-of-the-art methods struggle to achieve high accuracy in this area. This highlights a clear need for improved approaches that better handle legal reasoning.
@@
Legal NER focuses on specific challenges such as disambiguating titles, resolving nested entities, addressing co-references, managing lengthy texts, and processing machine-inaccessible PDFs. Despite its critical role in understanding and structuring legal documents, there is limited research in this area, as observed from Fig. 1.
@@
The concept of “Stochastic Parrots” pertains particularly to LLMs. It shows the concern that these models often do not truly understand language but merely mimic human patterns. This mimicry can lead to unreliable outcomes, especially in critical legal situations, if the models are not trained on high-quality, unbiased datasets. The risk is notably significant in LJP, where training on biased or unfair data could lead to irreversible outcomes, as discussed in Bender et al. (2021)’s work on the limitations of LLMs. This underscores the importance of ensuring that LLMs are trained responsibly to avoid perpetuating or amplifying existing biases in legal decisions.
@@
In the legal domain, where documents are usually lengthy, often contain cross-references, and present a variety of complex linguistic structures, LLMs can sometimes generate hallucinatory responses when faced with the task of generating accurate answers. RAG systems offer a promising solution to these challenges. RAG can mitigate issues such as the natural limitations of LLMs concerning maximum input lengths, where even extended limits may fall short due to the excessive length of many legal documents. This approach not only improves the model’s response quality but also its relevance and contextual appropriateness by incorporating more of the document’s content into the decision-making process. However, the integration of RAG into the legal domain introduces unique challenges, such as managing documents from multiple jurisdictions, ensuring temporal relevance, addressing multilingual issues, and overcoming biases in the retrieval phase. These challenges must be addressed in future research when applying RAG in the legal domain.
@@
Table 4 illustrates the connections between ORCs and discussed areas. A direct relationship is marked with an ✓, and otherwise with a –. As shown, most ORCs are related to LJP, LQA, LTC, and LLMs, indicating more extensive research fields in these areas.
Table 4. Summary of existing ORCs in each area.
@@ 14. Conclusion
Advances in AI and NLP have improved legal NLP techniques and models. These improvements help better meet the needs of laypersons in legal matters and ease the workload of legal professionals. This survey provides a comprehensive overview of the advancements in NLP techniques used in the legal domain. Additionally, we discussed the unique characteristics of legal documents. We also reviewed existing datasets and LLMs tailored for the legal domain. Legal NER research spans multiple languages and utilises diverse methods, from rule-based to BERT-based models. LDS has largely focused on extractive and abstractive methods, including TF-IDF and transformer-based models. In LTC, multi-class classification tasks dominate, with deep learning architectures, such as CNNs and Bi-LSTMs widely used. LJP primarily focuses on Chinese datasets with deep learning approaches, such as CNNs. LQA often leverages IR techniques such as BM25, with a significant focus on statutory law. Finally, we explored key ORCs, such as the need for domain-specific fine-tuning strategies, addressing bias and fairness in legal datasets, and the importance of interpretability and explainability. Other challenges include the development of more robust pre-processing techniques, handling multilingual capabilities, and integrating ontology-based methods for more accurate legal reasoning.

### v3

@@ 13. Open Research Challenges
Pre-processing legal texts is challenging because legal documents often consist of raw texts that requires extensive cleaning and transformation before use in ML models. Their complex, nested structures, such as clauses within clauses and cross-references to other cases, statutes or provisions, make it hard to segment them into coherent units for analysis. Without addressing these complexities, fine-tuning LMs on raw legal data becomes impractical, limiting the performance of legal NLP applications.
@@
The use of RLHF within the legal domain is notably scarce. Currently, there is only one peer-reviewed work (Nguyen et al., 2021) available that explores this approach. This indicates an opportunity for research and development in this area, as RLHF could potentially enhance the capability of NLP models to learn and to make decisions based on complex legal data under human guidance. Further exploration into this method could lead to more responsive and adaptable legal NLP systems. However, due to the complex nature of legal reasoning and the need for accurate legal knowledge in the human feedback phase, integrating RLHF into legal NLP pipelines poses some challenges. In particular, legal experts such as lawyers and judges must provide guidance to ensure the AI models accurately interpret and apply complex legal concepts, which is a barrier due to the significant cost of employing such professionals. Nonetheless, if RLHF can demonstrate further downstream savings due to improved worker efficiency, the up-front cost of employing legal professionals for data annotation tasks may be a worthwhile investment.
@@
There is a noticeable gap in the research across various areas of the legal domain, including Intellectual Property, Criminal Law, Banking Law, Family Law, and Human Rights Law. These fields have seen limited exploration across all legal NLP tasks, such as LQA and other applications. Expanding research into these areas is essential for developing comprehensive automated legal systems that can provide tailored solutions and insights highly relevant to these sectors of law.
@@
Research into SLMs specific to the legal domain is notably absent. Although some work explored notions of distillation and the use of smaller transformers, there has not been a significant exploration into the efficiency considerations of employing large language models in the legal domain. Addressing this gap could lead to more efficient, resource-conscious solutions that still maintain high performance in legal text processing and analysis. The development of SLMs tailored for legal applications could revolutionise the accessibility and scalability of legal NLP tools.
@@
Domain-specific efficient fine-tuning within the legal field remains underexplored, with only two known studies addressing it (Louis et al., 2024; Li et al., 2022). Legal texts feature complex structures and specialised vocabulary that standard LLMs may not capture without substantial adaptation. Moreover, the legal domain covers a vast array of document types, such as case law, statutes and contracts, each requiring tailored model strategies. This diversity makes it imperative to develop fine-tuning methods that not only adapt a model generally but also tailor it to the specific characteristics of each document type. Most existing approaches fine-tune the entire model, which can be resource intensive. More focused research could enable efficient fine-tuning of legal LLMs using fewer resources and improve their practical deployment.
@@
Complex legal logical reasoning remains a challenge in LJP, particularly in predicting prison terms. Current SOTA methods struggle to achieve high accuracy in this area, highlighting a clear need for improved approaches that better handle legal reasoning. Multi-hop QA Models with knowledge graphs, or reasoning LMs, may be an avenue worth investigating here.
@@
Legal NER focuses on specific challenges such as disambiguating titles, resolving nested entities, handling co-references and processing PDFs that are not machine-readable. Despite its crucial role in structuring and interpreting legal documents, research on legal NER remains limited, as shown in Figure 1.
@@
The concept of “Stochastic Parrots” pertains particularly to LLMs. It shows the concern that these models often do not truly understand language but merely mimic human patterns. This mimicry can lead to unreliable outcomes, especially in critical legal situations, if the models are not trained on high-quality, unbiased datasets. The risk is notably significant in LJP, where training on biased or unfair data could lead to irreversible outcomes, as discussed in Bender et al. (2021)’s work on the limitations of LLMs. Ultimately, we must ensure that tools and algorithms aimed at automating legal processes do not become decision makers, rather that they are framed as decision support tools. More research is required to fully grasp the sociotechnical aspects of NLP in the legal domain.
@@
The legal domain presents particular challenges because documents are lengthy, contain numerous cross-references, and exhibit complex linguistic structures. These characteristics can cause LLMs to hallucinate when asked for precise answers. RAG offers a promising remedy by incorporating relevant passages directly into the generation process, overcoming LLM input-length limits and improving both relevance and contextual accuracy. However, applying RAG in legal settings still poses challenges: handling documents from multiple jurisdictions, ensuring that retrieved material remains temporally current, addressing multilingual texts, and mitigating retrieval biases are all significant problems that must be resolved.
@@
Table 5 illustrates the connections between ORCs and the discussed research areas. As shown, most ORCs are related to LJP, LQA, LTC and LLMs, indicating more extensive research in these areas. Nonetheless, it is evident that there remains significant work to be done in all areas of focus, and this work is likely to be of a cross-disciplinary nature, spanning a number of research communities.
Table 5. Summary of existing ORCs in each area. A direct relationship is denoted with a check-mark (✓).
@@ 14. Conclusion
Advances in AI and NLP have improved legal NLP techniques and models, reducing the difficulty of engaging in legal processes for laypersons, and easing workloads and manual labour for professionals. This survey provides a comprehensive overview of the advancements in NLP techniques used in the legal domain, paying special attention to the unique characteristics of legal documents. We also reviewed existing datasets and LLMs tailored for the legal domain. Legal NER research spans multiple languages and utilises diverse methods, from rule-based to BERT-based models. LDS has largely focused on extractive and abstractive methods, ranging from TF-IDF to transformer-based models. LAM now automates the detection of claims, premises, and their links through domain-specific annotation schemes and graph-based or residual-network approaches, supporting legal reasoning tasks such as conflict resolution. In LTC, multi-class classification dominates, with DL architectures, such as CNNs and Bi-LSTMs widely used. LJP primarily focuses on Chinese datasets with DL approaches, such as CNNs. LQA often leverages IR techniques such as BM25, with a significant focus on statutory law. Finally, we explored key ORCs, such as the need for domain-specific fine-tuning strategies, addressing bias and fairness in legal datasets and the importance of interpretability and explainability. Other challenges include the development of more robust pre-processing techniques, handling multilingual capabilities and integrating ontology-based methods for more accurate legal reasoning.